At the Golden Oldies Theater,tickets for adults cost $5.50 and tickets for children at $3.50. How many of each kind of ticker were purchased if 21 tickets were bought for $83.50?
Found 2 solutions by harpazo, mananth:
Answer by harpazo(655)

This is a system of linear equations problem that you can solve using the substitution method. I will set up both equations and you do the math.
Let a = adult tickets
Let c = children tickets
a + c = 21
5.50a + 3.50c = 83.50


Answer by mananth(16946)

x= Number of adult tickets
y= Number of child tickets

1.00 x + 1.00 y = 21.00 .............1
Total value
5.50 x + 3.50 y = 83.50 .............2
Eliminate y
multiply (1)by -3.50
Multiply (2) by 1.00
-3.50 x -3.50 y = -73.50
5.50 x + 3.50 y = 83.50
Add the two equations
2.00 x = 10.00
/ 2.00
x = 5.00
plug value of x in (1)
1.00 x + 1.00 y = 21.00
5.00 + y = 21.00
y = 21.00 -5.00
y = 16.00
y = 16.00
x= 5 Number of adult tickets
y= 16 Number of child tickets